To appear in Fancy J. Version of 15 Sep. 2004

Article Title

by You and Me

Abstract. This is a template file for articles to illustrate the use of various macro packages for plain TEX.

Contents

1. Introduction 1

2. Background 2 References 2

§1. Introduction.

Here are some bibliographic references: Joag-Dev and Proschan (1983) and Pemantle (2000).

Here are aligned equations:

P^{f}[η(e_{1}) = 1, …, η(e_{k}) = 1] := P^{f}[{η ∈ {0, 1}^{ℤ^{d}} ; η(e_{1}) = 1, …, η(e_{k}) = 1}]
:= det[f̂(e_{j} − e_{i})]_{1≤i,j≤k} .  (1.1)

We give a forward reference to Section 2.

Example 1.1. Here’s an example.

2000 Mathematics Subject Classification. Primary 28D05. Secondary 37A60.

Key words and phrases. Macros.

Research partially supported by money.

§2. Background.

Here’s a theorem.

Theorem 2.1. If Q is any positive contraction on ℓ^{2}(E), A is a finite subset of E, and η_{0} ∈ 2^{A}, then P^{Q}[ · | η↾A = η_{0}] has negative associations.

Acknowledgements. Thanks for reading this.

REFERENCES

Joag-Dev, K. and Proschan, F. (1983). Negative association of random variables, with

applications. Ann. Statist. 11, 286–295. Pemantle, R. (2000). Towards a theory of negative dependence. J. Math. Phys. 41,

1371–1390. Probabilistic techniques in equilibrium and nonequilibrium statistical physics.

Department of Mathematics, Indiana University, Bloomington, IN 47405-5701

rdlyons@indiana.edu
http://mypage.iu.edu/~rdlyons/